# ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion

### v1

@@ 3.2 Model Collapse in the Chain of Diffusion
CFG scale does not only affect the speed of model collapse, but also the pattern of model collapse. As shown in Figure 22(b), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to low-frequency degradation. On the other hand, for CFG $7.5$ how images degrade looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated. The t-SNE plot in Figure 33(a) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales. These patterns were consistent in all four datasets and detailed results can be found in Appendix B.
We further analyze this phenomenon by examining the latent space. First, we show a histogram of the final latent vectors before decoding in Figure 44(a). For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations. This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode. We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space. For a CFG of $7.5$, the evolution of the latent distribution is completely opposite, creating longer tails and approaching a more uniform distribution over space. For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations.
@@
Implications of our observations.
Through extensive investigations, we gained a comprehensive understanding of CFG scales’ effects on the Chain of Diffusion. Even though a high CFG scale of $7.5$ is a common choice to generate visually appealing images, it significantly sacrifices reusability to achieve slightly better FID. Moreover, it is noteworthy that CLIP score increases for the early three iterations and drops rapidly in the latter when the CFG scale is $7.5$ (Figure 33(b)). This shows that Stable Diffusion tends to generate images favored by CLIP encoder, which can contain artifacts easily amplified by further finetuning. Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [12] and they provide FID plots with varying levels of sampling bias by changing the CFG scale. While they reported a monotonic decrease in reusability as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$. It shows an intriguing trade-off between perceptual quality and reusability. This suggests that if developers of diffusion models are mindful of the reusability metric, can improve future generations of images substantially by carefully choosing CFG.
4 ReDiFine
In the previous section, we showed that choosing a good CFG scale can considerably improve reusability in the Chain of Diffusion. However, such optimal CFG scales vary from dataset to dataset. To mitigate model collapse solely with CFG tuning, we have to do multiple rounds of finetuning and sampling to evaluate the reusability-FID trade-off. This is too computationally burdensome to be a practical solution. In this section, we propose ReDiFine, Reusable Diffusion Finetuning, which can mitigate model collapse without any parameter tuning. ReDiFine combines two ideas—condition drop finetuning and CFG scheduling—and effectively improves the reusability-FID trade-off in all four datasets (comparable to optimal CFG scales).
4.1 Condition drop finetuning.
We propose condition drop finetuning, which randomly drops the text condition during finetuning to update both conditional and unconditional scores. Even though condition drop was suggested in the original CFG paper [32], it is not commonly done during finetuning since the first-iteration image qualities are outstanding without updating unconditional scores. However, the small difference between conditional and unconditional scores can accumulate over iterations and create a large gap as the iteration progresses in the Chain of Diffusion if we do not explicitly finetune the unconditional scores as shown in Figure 44(b). Inspired by this observation, we drop text conditions with probability $0.2$ during finetuning. We illustrate the effect of condition drop with CFG $7.5$ in Figure 44(b). Note that the norm of Diff does not grow as fast even with a high CFG of $7.5$. While this prevents a potential cause of the high-frequency degradation, the Diff value still spikes up in iteration 6. We conjecture that this is because the $D_{5}$ already contains severe artifacts that conditional scores have to grow substantially to fit the data.
4.2 CFG Scheduling.
Based on the analysis showing that a high CFG scale amplifies the discrepancy between the condition and uncondition scores during iterative fine-tuning (see Figure 44(b)), we propose gradually reducing the CFG scale during diffusion steps. This CFG scheduling mechanism mitigates the negative impacts of overemphasizing the conditional score in later stages while preserving the initial conditional information in the early stages. Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as
$s=s_{0}\times e^{-\alpha\times t/T}$ (6)
where $T$ is $30$ and $s_{0}$ is $7.5$. $\alpha$ denotes the degree of decrease and we use $2$ throughout this paper. Comparison to different $\alpha$ and linear decreasing can be found in Appendix E.
Our method gradually decreases the CFG scale during diffusion steps, leveraging the benefits of high CFG scale in the early stage and low CFG scale in the later stage. This adaptive approach aligns with observations from [34] that different steps contribute uniquely to the generation process. In addition, our CFG scheduling takes advantage of using a small CFG scale without the need for extensive hyperparameter searches for each dataset, providing a flexible and robust choice for the Chain of Diffusion.
5 Reusable Image Generation
@@
Figure 6: ReDiFine effectively mitigates Reusability-FID trade-off across all four datasets. While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low reusability and FID at the same time (lower is better). While CFG scheduling alone works well for Pokemon dataset, it is still beneficial to have condition drop finetuning for CelebA-1k and Kumapi datasets. Note that the differences in FID are relatively smaller than those in reusability, supporting the necessity to evaluate reusability in the Chain of Diffusion.
In this section, we present generated images and qualitative metrics for ReDiFine. We use the CFG scale $s_{0}=7.5$ for all datasets and experiments except for the ablation study. Note that ReDiFine can be further improved with hyperparameter tuning, e.g., CFG scale, condition drop probability, and CFG scheduling function. However, our focus in this work is to demonstrate that ReDiFine can robustly stabilize iterative finetuning regardless of datasets and hyperparameters. This robustness provides a great advantage when applying our ReDiFine to finetune Stable Diffusion and generate images using any personally collected image sets.
Visual comparison.
Comparing Figure 5 with Figure 11(b), ReDiFine effectively slows down and mitigates the image degradation in the Chain of Diffusion compared to using the default CFG scale $7.5$ across four datasets. In addition, ReDiFine successfully achieves the same level of high-quality image generation in the Chain of Diffusion compared to using the optimal CFG scales for each dataset, which requires extensive hyperparameter tuning. This is demonstrated by visually comparing the generated images in Figure 5 with the ones generated using the optimal CFG scales for Pokemon and CelebA dataset in Figure 22(b). Additional comparisons for Kumapi and Butterfly datasets can be found in Appendix B.
Mitigate Reusability-FID trade-off.
In addition to the visual comparison, we also provide quantitative results in Figure 6. We can see that ReDiFine consistently helps mitigate Reusability-FID trade-off in the Chain of Diffusion across all four datasets. In contrast, a fixed CFG scale that works best on one dataset fails on others. For example, the optimal CFG scale $2.5$ for Pokemon dataset achieves high (poor) ${\rm Reusability}$ for CelebA-1k and Butterfly datasets. Moreover, the optimal CFG scale $1.5$ for CelebA-1k and Butterfly datasets results in the high (poor) ${\rm FID}_{1}$ for Pokemon dataset. In addition to its ability to generalize across different datasets, ReDiFine demonstrates robustness when initialized with different CFG scales. Illustrated in Figure 66(a), ReDiFine achieves comparable performance with CFG scales of $5.0$ and $10.0$ as CFG $=7.5$.
Ablation study.
We provide an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine. As shown in Figure 6, CFG scheduling alone achieves comparable reusability in the Pokemon dataset, but condition drop finetuning is essential for CelebA-1k and Kumapi. Combining these two, ReDiFine consistently achieves the best trade-off between ${\rm Reusability}$ and ${\rm FID}_{1}$ across all four datasets. Additionally, the differences in ${\rm FID}_{1}$ are relatively small—using different CFG scales can generate high-quality images in the first iteration except for very low CFG scales, e.g., CFG = $1.0$. In contrast, the variance of ${\rm Reusability}$ is much larger. This supports the necessity of evaluating our proposed ${\rm Reusability}$ metric for the Chain of Diffusion. Additional results can be found in Appendix E.
More iterations.
In addition to the comparison over 6 iterations, Figure 7 illustrates the evolution of image degradation over 12 iterations. Notably, as fine-tuning iterations increase, high-frequency degradation emerges with the optimal CFG, highlighting the necessity for additional hyperparameter search to sustain image quality across longer diffusion lengths. Conversely, ReDiFine preserves image quality without high-frequency degradation.
Figure 7: Comparison of the optimal CFG scale and ReDiFine on 12 iterations for Pokemon dataset.
(a) Histogram of latent values.
(b) Norm of differences.
(c) Power spectrum of latents.
Figure 8: Comparison of latent analysis at the 6th iteration for ReDiFine and baselines. (a) ReDiFine preserves the latent distribution. (b) Conditional and unconditional score differences do not surge in the $6$-th iteration, showing dramatic improvements from baselines. (c) ReDiFine maintains similar power spectrum to CFG scale $2.5$ in the $6$-th iteration.
Latent analysis.
Figure 8 demonstrates that ReDiFine can successfully alleviate the symptoms we observed in the latent space in Section 3. When we plot latent histograms, we can notice that ReDiFine preserves the original latent distribution at the 6th iteration as compared to other baselines with different CFG values. More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores. Additionally, the power spectrum of latents in the $6$th iteration increases or decreases for CFG scales $1.0$ and $7.5$ while ReDiFine and CFG scale $2.5$ preserve density distribution. More analysis and comparison for all iterations are provided in Appendix F.1.
(a) Autocorrelation.
(b) Average power spectra.
Frequency fingerprint.
Several works [35, 36] aim to identify fingerprints of synthetic images. High-quality synthetic images from different generative models have clearly distinct fingerprints, showing the potential to be used for synthetic image detection. We analyze fingerprints of synthetic images for different CFG scales and iterations, and compare them to fingerprints of the original training set. Both power spectra and autocorrelation show clear differences between the original training set and synthetic images as shown in Figure 9. Moreover, fingerprints of synthetic images from ReDiFine are similar to those of images from CFG scale $2.5$, while other CFG scales ($1.0$ and $7.5$) make fingerprints different from the first iteration as iterations proceed. Horizontal and vertical lines in autocorrelation gradually disappear and central regions in power spectra vary for further iterations. Also, the varying central regions in power spectra imply that low frequency features increase and decrease for CFG scale $1.0$ and $7.5$, respectively, aligning with visual inspections. Generating images with fingerprints similar to those of the original real images can be an interesting future direction to reduce the degradation in the Chain of Diffusion. Additional analysis with power spectrum density of images is provided in Appendix F.2.
6 Conclusion and Discussion
Motivated by the prominent usage of diffusion models by artists and the trend of the Internet being filled with synthetic data, we investigate how image quality degrades as we feed diffusion-generated synthetic images to finetune another diffusion model. In our self-consuming chain of diffusion finetuning setting, we observe image degradation in all four datasets we tested. We reveal that CFG plays a crucial role in the trade-off between image reusability and its perceptual quality in the first iteration. We propose a novel idea of generating reusable iamges and propose the ReDiFine strategy that combines two small techniques (condition drop finetuning and CFG scheduling) that can be easily implemented in the diffusion finetuning pipeline with no requirement for additional parameter tuning.
We started this paper with a question: can current AI models learn from their own output and improve themselves? This is not only an imminent question as synthetic data rapidly outnumbers human-generated data [37, 38], but also a long-term existential question as it relates to reaching AI singularity—the point where AI learns from its own output to continuously enhance its intelligence and far surpasses human intelligence. Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet. However, we believe that there are many open questions on how we develop AI in a world dominated by AI data. We present one angle by emphasizing the importance of generating reusable data for future AI developers. Another pathway could be developing a new learning algorithm that can discern real and synthetic data and take a different training technique for each.

### v2

@@ 3.2 Model Collapse in the Chain of Diffusion
The CFG scale affects affect not only the speed of model collapse, but also the pattern of model collapse. As shown in Figure 33(a), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to as low-frequency degradation. On the other hand, for CFG $7.5$ the degradation pattern looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated. The t-SNE plot in Figure 33(b) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales. These patterns were consistent in all four datasets and detailed results can be found in Appendix B. In Section 4, we detail how different patterns of power spectra in high-frequency regions from different CFG scales can be understood using the framework from genetic biology.
Implications of our observations. Our extensive investigations reveal that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$. Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [16]. While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$. It shows an intriguing trade-off between perceptual quality and reusability. This suggests that developers concerning reusability of images can substantially improve future generations by carefully choosing CFG.
4 Quantitative Trait Models for Fully-Synthetic Training Loops
This section introduces a novel perspective for understanding the model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion. Distinguishable iterations in the Chain of Diffusion—where each iteration is clearly separable with no duplicated individuals, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring. By applying quantitative trait modeling from statistical genetics, we provide a framework to describe how images evolve across iterations in the Chain of Diffusion.
We begin by introducing quantitative trait modeling and its underlying mathematical assumptions. Leveraging these assumptions, we derive a theorem demonstrating that the mean trait value exhibits linear drift while the variance stabilizes over time. We then show that this model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale performance. This suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits. We use the term “iteration” instead of “generation” for genetic generation to avoid confusion with image generation.
(a) Selection mechanism.
(b) Power spectra of images vs. Mean of phenotypes.
Figure 4: (a) Selection mechanism with two-sided truncation. $r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected. (b) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results of different selection strategies. Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale. Detailed settings are provided in Appendix D.2.
4.1 Quantitative Trait Modeling
Quantitative trait modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are determined by multiple genetic and environmental factors in complex ways. These phenotypes are typically assumed to follow a Gaussian distribution, with discrete iterations where parents and offspring are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
Let the distribution of phenotypes at the $t$-th iteration be denoted as $\mathcal{N}(\mu_{t},\sigma_{P,t}^{2})$ where $\mu_{t}$ and $\sigma_{P,t}^{2}$ are the mean and variance of quantitative phenotypes. The phenotypic variance $\sigma_{P,t}^{2}$ is the sum of the (additive) genetic variance $\sigma_{G,t}^{2}$ and the environmental variance $\sigma_{E}^{2}$, i.e., $\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}$ [38] [^6] Genetic variance is composed of additive, dominance, and interaction variance. Here, we only consider additive variance, which is a common assumption in the field.. When selection occurs in each iteration, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration. We consider directional selection with truncation, as shown in Figure 44(a), where a ratio of $r$ samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution. Here, $r+r_{1}+r_{2}=1$ and larger phenotype values are preferred when $r_{1}>r_{2}$.
The (narrow-sense) heritability [^7] The heritability is narrow-sense when the genetic variance is restricted to additive variance., which is defined as the proportion of phenotypic variance attributable to additive genetic factors, can also be represented using the Breeder’s Eqn. [39] as:
$h_{t}^{2}=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}}=\frac{\sigma_{G,t}^{2}}{\sigma_{G,t}^{2}+\sigma_{E}^{2}}=\frac{\mu_{t+1}-\mu_{t}}{\mu_{t}^{\prime}-\mu_{t}},$ (5)
where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population at current iteration $\mu_{t}$, and heritability $h_{t}^{2}$. Additionally, we assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$. We prove the behaviors of mean and variance of phenotypes under this setting:
Theorem 1
Suppose the distributions of phenotypes follow Gaussian distribution and directional selection truncates individuals on both sides with ratios $r_{1}$ and $r_{2}$. Then mean of phenotypes asymptotically increases (decreases) by $\frac{c_{1}c_{2}}{\sqrt{1-c_{2}}}\sigma_{E}$ per iteration and the variance converges to $\frac{1}{1-c_{2}}\sigma_{E}^{2}$ when $r_{1}>r_{2}$ ($r_{2}<r_{1}$), where $c_{1}$ and $c_{2}$ are constants that depend on $r_{1}$ and $r_{2}$.
The proof of Theorem 1 is provided in Appendix D.1.
4.2 Explaining the Chain of Diffusion with Quantitative Trait Modeling
Just as phenotypes evolve through complex functions of hidden genotypes across multiple iterations in quantitative trait modeling, various continuous features of images evolve similarly during the Chain of Diffusion. A key discovery in our work is that the high-frequency power spectra are crucial phenotypes that evolve over iterations, with the CFG scales acting as different selection mechanisms. These power spectra are computed using 2D Fourier transforms, focusing on frequencies above a certain threshold to capture high-frequency components. Different CFG scales during generation correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
Figure 44(b) shows the evolution of power spectra in the Chain of Diffusion (in solid line) and the phenotype mean modeled with Equation  (9) (in dotted lines). The initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$ in the simulation are set to match those of the original image set at iteration 00. Three different configurations of ratios, $r_{1}$ and $r_{2}$, successfully model the power spectra distribution: CFG $7.5$ is modeled as selecting from upper $5$% of samples, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ results from selecting the lower $30$%. We can see that this modeling can capture the evolution of power spectra very accurately over 6 iterations of Chain of Diffuson. Details about power spectra computation and simulation parameters are provided in Appendix D.2.2.
4.3 Statistical Genetics as a Unifying Lens to Model Collapse
Quantitative trait modeling, a tool borrowed from statistical genetics, serves as a solid theoretical framework for interpreting our experimental results. As genetic processes and the self-consuming training of generative models share a lot of similarities, we find that other concepts in mathematical genetics also have close ties to existing theoretical work on model collapse. For instance, the traditional Wright-Fisher model [19] describes how traits evolve from one generation to the next in a finite population. Due to randomness in the sampling process, the composition of traits within a finite population drifts over time and eventually collapses to a single phenotype. The Wright-Fisher model and its continuous variants [40] are exactly equivalent to the collapse behavior modeled with simple categorical or Gaussian distributions in [16, 3, 6, 22]. Given its widespread use in genetics, many extensions of the Wright-Fisher model remain an active area of research, such as including different mechanisms of selection [41, 42] or mutations [43]. Beyond the current work and existing theoretical studies on model collapse, we believe that statistical genetics offers a unified perspective to understand model collapse by reducing the complex dynamics of the self-consuming loop of generative model training to a small number of parameters, from which we can gain valuable insights.
5 Reusable Image Generation with ReDiFine
In the previous sections, we have discovered a significant role of CFG in model collapse and that an appropriate choice of CFG can effectively mitigate the collapse. However, the optimal CFG value that minimizes the collapse rate is different for each dataset (e.g., CFG 1.5 for CelebA, CFG 2.5 for Pokemon). There is no efficient way of finding an optimal CFG other than performing a grid search over a wide range of values and iteratively fine-tuning the diffusion model to evaluate the collapse rate for each configuration. In practical scenarios, it is unlikely that end users who are not machine learning experts will go through such a process during finetuning, just to prevent a potential model collapse that can happen in the next generation. This raises the question: How can we design a finetuning and generation strategy that is user-friendly and can effectively slow down model collapse?
To address this question, we again draw inspiration from the evolution process in nature. In biology, mutations naturally counteract genetic drift and preserve diversity. Furthermore, selection in nature is often a soft process rather than a hard truncation, as illustrated in Figure 44(a). This soft selection allows for the inclusion of outliers, thereby maintaining the overall genetic diversity. We connect these biological inspirations with two strategies: condition drop finetuning to introduce more randomness and CFG scheduling during generation to transform hard selection to a softer one [^8] We show a modified quantitative trait modeling result with these two modifications in Appendix D.2 and show that it captures the ReDiFine experimental results effectively. In this section, we focus on showing the image generation results with ReDiFine. We refer the curious readers to the appendix.. We propose Reusable Diffusion Finetuning (ReDiFine) method that incorporates these two ideas and show that it is highly effective at mitigating model collapse without the need for any finetuning.
@@ 5 Reusable Image Generation with ReDiFine
CFG Scheduling. We propose gradually reducing the CFG scale during diffusion steps to mitigate the impacts of overemphasizing the conditional score in later stages, which can lead to high-frequency degradation. Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as $s=s_{0}\times e^{-\alpha\times t/T}$, where $s_{0}$ is the initial CFG scale and $\alpha$ is the rate of exponential decay. This scheduling approach is consistent with findings by [44], which suggest that different diffusion steps contribute uniquely to the generation process. High CFG lets us capture high-level semantic information accurately during the initial diffusion steps while lower CFG in later steps can prevent generating unnecessary high-frequency details.
@@
Figure 6: ReDiFine effectively mitigates collapse-FID trade-off across all four datasets. While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better). Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
ReDeFine Results. We present the generated images and quantitative metrics for ReDiFine. We use the initial CFG scale $s_{0}=7.5$, diffusion steps $T=30$, decay rate $\alpha=2$, and condition drop probability $0.2$ for all of our experiments, except in the robustness comparison and ablation study. While ReDiFine can be further optimized through hyperparameter tuning (e.g., changing condition drop probability), our goal is to demonstrate that ReDiFine robustly mitigates model collapse with these default parameter settings for all datasets. In Figure 5, we can clearly see that ReDiFine significantly improves the image quality at later iterations compared to the baseline (Figure 11(b)) for all four datasets. In addition to the visual comparison, Figure 6 quantitatively shows the collapse-FID trade-off of ReDiFine. In all four datasets, ReDiFine shows substantially lower collapse rate (y-axis), compared to the baseline case (CFG=7.5). Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values across all four datasets, demonstrating its effectiveness as a universal solution for mitigating model collapse. In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA but performs poorly for Pokemon and Kumapi.
We also compare ReDiFine with real image mixing, which is suggested as a mitigation strategy in several previous papers [16, 3, 8, 23]. In Figure 7, we can observe that even infusing the dataset with 90% real images, it shows severe image degradation within six iterations. By visual comparison, we can see that Pikachu images in Figure 5 have less degradation than those with real image mixing. Furthermore, applying ReDiFine to these mixed datasets, as shown in Figure 7, significantly improves image quality compared to the baseline. This demonstrates that ReDiFine is an effective solution for various settings and can be used alongside other mitigation strategies such as real data mixing.
Figure 7: Comparison of ReDiFine and baseline (CFG scale 7.5) with training sets mixed with original and synthetic images from the previous iteration. While mixing real images as much as $90$% still results in model collapse for the baseline (as shown in the first row), the ability of ReDiFine to mitigate model collapse is even more highlighted by mixing real images (second row).
Ablation study & Further analyses. We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine. We plot the results of using only condition drop and CFG scheduling in green triangles on Figure 6. In Pokemon, using only one of those strategies outperforms ReDiFine, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine. Especially in CelebA, using either one of them showed significantly worse performance than ReDiFine. These results suggest that combining both condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets. We conducted further analyses on ReDiFine, examining the distribution of latent features, the evolution of the norm of Diff, the power spectra using 2D Fourier transforms, and forensic fingerprints based on prior studies [45, 46]. Our analysis shows that ReDiFine effectively preserves the latent distribution and the norms of Diff over six iterations, with forensic fingerprints closely resembling those of the optimal CFG case. Detailed results are provided in Appendix G.
Reusable image generation as a responsible AI practice. In this section, we demonstrated that ReDiFine significantly slows the collapse rate without sacrificing image quality, similar to how fair classifiers reduce bias without compromising accuracy [47]. Adopting ReDiFine is a responsible, environment-conscious practice to prevent polluting the world with images that might look visually appealing, but totally unusable to improve future AI models.
6 Conclusion
The influx of AI-generated data into the world is inevitable and training sets that consist of synthetic data will be part of the AI development pipeline. In this paper, we empirically and theoretically studied the scenario of finetuning a model with its own generated data, where a gradual degradation called “model collapse” happens. We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose an effective mitigation strategy that generates reusable images for future training.
We started this paper with a question: can current AI models learn from their own output and improve themselves? Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet. While we presented one solution focused on generating reusable data, many open directions remain, such as developing algorithms that can distinguish between real and synthetic data and apply different learning techniques accordingly.